Contingency Engineering Plan for UT Nov 07-09, 2019

  • DONE Testing whether WFS_probeRotation could be set to the wrong sign for AGW1 . 1 hour - any conditions - Highest Priority for non-ARGOS Eng - J. Hill
    (This is a possible solution to IT 5142 - causing active optics runaways on particular standard fields.)
    • Authorize and configure as LUCI/LUCI (LUCI2 is optional)
    • Send a binocular ACTIVE preset with the off-axis guide star far out on the X-axis (X at least +-140 mm, preferably 145 mm). Try to avoid the vignetting of IT 7235, so one of +-140 mm might be preferred.
      • Target can be anything with a suitable guide star - we just want to see if active optics runs away or not.
      • Check guide star position with 'gcsclient getStatus left'
    • Let active optics run for a few minutes until it starts to runaway in Z9/Z10 (Steve is expert in detecting this.) The experiment is still useful if we don't see a runaway here.
    • Change the sign of WFS_probeRotation in LUCI_L.cfg from 1.0 to -1.0 (SA: Do you mean WFS_probeRotationFactor?)
      • as telescope@obs1, 'ssh tcs@tcs1'
      • on tcs1, cd /lbt/data/config/tcs/GCS/
      • vi LUCI_L.cfg
      • stop and restart GCSL
    • Resend same binocular ACTIVE preset with the off-axis guide star far out on the X-axis
    • Active optics stability should either be better, or much worse. Keep or revert the change as needed.
  • LUCI pseudo-speckle observations - 1 hour - D. Thompson
    • SL on LUCI-1, ESM/DL-AO on LUCI-2
    • 3 sources availble (observe only 1)
    • Requires GEIRS with SRR mode, cube-saved mode, sub-array readouts
  • Photometric Standards, LUCI, MODS, LBC - any seeing - clear skies
    • 20-25 min per target minimum 2 visits per instrument over different airmasses
  • DONE Test 2019B TCS Build - 10 min each way to setup/revert, for bad weather
    • To be installed if ARGOS is not running
  • DONE Test PCS patch for IT 7640 (problems with LUCI/SOUL repoint=yes acquisitions) can be seeing up to 2 arcsec
    • send ACE_ACTIVE preset, do acquireref with repoint=yes; PCS patch will be activated in TCS 2019A
  • Continue to collect ESM data - all ARGOS targets have SL+ESM and ESM-LUCI-2 versions of scripts
    • M52 - 45 min per AO Ref star script - optimal is to do both (1.5 hours) - sets 2:30am
    • HCG 31 (if time) - if only one M52 script can be done, then do HCG 31 script (45 min) - up all 2nd half
    • M34 - 45 min per script - sets 5am
    • M103 - 45 min per script - sets 4:30am
    • DONE M67 - 45 min per script - 2am onwards
  • If poor conditions (poor seeing and/or thick clouds) - backup science:
    • H-band spectral library - Bino
    • Arp 187 LS - Bino
  • LBC superfoc sequences - requires ~1" stable seeing and can be LBCR only (contingent on results for blue which are coming)
    • lbto@obsN:~/Startup/Startup_Sep_2019/LBC
    • need a dense star field - Possible fields are those for the copointing OBs in the "Sci-like OBs" group in the OT program file LBC_Checkout_v1.6.xml.
    • Sequence - repeat several times
      • dofpia
      • superfoc at PA=0

      • superfoc at PA=180
  • LUCI + LBC mixed mode --- any conditions (more run time, in particular with and without dithering)
Scripts for LUCI to be placed in:
  • /home/observer/share/Service_Observing/Eng_Nov_2019 (RM507-4 and RM507-5)
  • Scripts for LBC and MODS (unlikely to be used): /home/lbto/Service_Observing/E_tasks

UT 07-Nov-2019

UT 11:50 - giving up on ARGOS - reconfiguring for LUCI-Bino, will run M67 SX-SL DX-ESM

UT 12:04 - collimation not great - moving to M67

UT 12:11 - loops closed, script starting - but not starting exposures

UT 12:12 - not sure why it hung, did skip to line and go - now taking exposures

UT 12:15 - co-pointing error even though it was a sync offset to a sky frame

ESM showing ~0.45" , SL showing ~ 1"

-factor ~2x improvement at J-band for ESM

L1 0028-0037 - K-band ; 0038-0047 J-band ; 0048-0057 H-band

L2 0003-0012 - K-band ; 0013-0022 J-band ; 0023-0032 H-band

UT 13:04 - finished M67 script - closing enclosure - blinds in flexure off

UT 13:15 - JHK imaging flats N375

L1 0058-0067 - J-band ; 0068-0077 - H-band ; 0078-0087 - K-band

L2 0033-0042 - J-band ; 0043-0052 - H-band ; 0053-0062 - K-band

UT 13:39 - cals complete - putting LUCIs in safe config


-- BarryRothberg - 29 Oct 2019
Topic revision: r17 - 09 Nov 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback